OpenStudy (anonymous):

hii can anyone help me with this question? prove that (sec x - tan x)(cosec x + 1) is equivalent to cot x

OpenStudy (cwrw238):

try converting LHS to sin's and cos's and simplify

OpenStudy (cwrw238):

might be a bit long-winded but this often works

OpenStudy (cwrw238):

i'll start you off i'll write c for cos x and s for sinx (cos i'm lazy)

OpenStudy (anonymous):

replace cosine by \(a\) and sine by \(b\) and multiply \[(\frac{1}{a}-\frac{b}{a})(\frac{1}{b}+1)\]

OpenStudy (cwrw238):

yea thats good idea

OpenStudy (anonymous):

\[(\frac{ 1 }{ cosx} - \frac{ sinx }{ cosx}) (\frac{ 1 }{ sinx } + 1)\] i'm not able to continue

OpenStudy (cwrw238):